Assume a model with no foreign sector, an income tax rate of 0.25, a consumption function defined as C = 680 + (0.8)YD and where autonomous investment, government spending and transfers are all equal to 100 and lump sum taxes are 200. What is the level of equilibrium income in this model economy?
1
Expert's answer
2017-02-21T13:47:11-0500
t=0.25
C = 680 + (0.8)Yd
I=G=T=100
TR=200
Equilibrium level of national income 
Y=C+I+G
Disposable income
Yd=(Y-T- t*Y+TR)
Therefore, substituting the values of various parameters and variables we have
Y = 680 + 0.8*(Y-T-t*Y+TR)+I+G
Y=680+0.8*(Y-200- 0.25*Y+100)+100+100
Y-0.8Y+0.2Y=680- 160+80+200
0.4Y=800
Y=2000
Answer: The equilibrium income in in this model economy is equal to 2000